A 150 lb., 62-year-old man had a chronic productive cough, exertional dyspnea, mild cyanosis, and marked slowing of forced expiration. His pulmonary function and laboratory tests follow:

Frequency

16 breaths/min

Alveolar ventilation

4.2 L/min

Vital capacity (VC)

2.2 L

Functional residual capacity (FRC)

4.0 L

Total lung capacity (TLC)

5.2 L

Maximum inspiratory flow rate

250 L/min

Maximum expiratory flow rate

20 L/min

PaO2

62 mm Hg

PaCO2

39 mm Hg

Pulmonary function tests after bronchodilator therapy:

Frequency

16 breaths/min

Alveolar ventilation

4.35 L/min

VC

2.4 L

FRC

4.0 L

TLC

5.2 L

Maximum inspiratory flow rate

250 L/min

Maximum expiratory flow rate

23 L/min

PaO2

62 mm Hg

PaCO2

38 mm Hg

  1. Calculate the minute ventilation (MV) for this person before and after the bronchodilator therapy.
  2. Is each MV normal or altered?
Answer #1

Minute ventilation(MV) is calculated by multiplying tidal volume( VT) by respiratory rate (RR).

MV = VT * RR

usually tidal volume is 600 ml in normal breathing person.

in this patient tidal volume before bronchodilator therapy is 422 approx. 400ml

so MV= 422* 16= 6752 ml or 6.7 L before bronchodilator therapy

tidal volume after bronchodilator therapy is 413 ml approx. 400 ml

So, MV = 413* 16= 6608 ml or 6.6 L after bronchodilator therapy

Minute ventilation MV is normal upto 9 L. more than 9L indicates hyperventilation.

in this patient MV is normal before and after bronchodilator therapy.